Assignment 2


1. In real-address mode, convert the following segment offset address to a linear address: 0950:0100


2. In real-address mode, convert the following segment offset address to a linear address: 0CD1:02E0


3. What is the duration of a single clock cycle (in nanoseconds) in a 3.4 GHz


4. A hard disk rotates at 4200 RPM (rotations per minute). What is the timeof one rotation in milliseconds?


5. Which Intel processor was the first member of the IA-32 family?

6. Which Intel processor first introduced superscalar execution?

7. Name all 32-bit general-purpose registers

8. Name all six CPU status flags in the Intel processor

9. Which flag is set when an arithmetic or logical instruction produces a
negative result?

10. What is the range of addressable memory in protected mode?
Answer #1

1. Given segment offset address(Segment:Offset pair) = 0950:0100

To find the linear address, append 0 to segment(0950h) and add 0100h to it.

It becomes 09500h+0100h = 09600h

2. Given segment address = 0CD1:02E0

To find the linear address, again append 0 to 0CD1 and then add 02E0 to it.

Linear address = 0CD10 + 02E0h = 0CFF0h

3.

Given frequency = 3.4GHz

1GHz = 109Hz

( 3.4GHz = 3.4*109 Hz )

Time duration of single clock cycle = 1 /(3.4*109 Hz )

= 0.294 * 10-9 Hz

= 0.294ns

Note : 1Hz = 1/ 10-9 ns

4.

Number of rotation in a minute = 4200

Time to take 1 rotation = 1/4200 min

= \frac{1*60000}{4200} msec

= 14.2msec

Note: 1 min = 60000millisec